Circular No. 8579

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 O2 (CHRISTENSEN) E. J. Christensen, Lunar and Planetary Laboratory, reports the dis-

covery of a comet (discovery observation below) in the course of the Siding Spring Survey, the new object showing a round coma ∼ 7′′ in diameter with a slight extension in p.a. ∼ 270^{o}, on images taken with the 0.5-m Uppsala Schmidt telescope.

2005 UT  α_{2000}  δ_{2000}  Mag.
July 31.75168  3^{h}34^{m}52.^{s}50  −11^{o}47′14′′.3  18.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-P21.

T = 2005 Sept.29.550 TT  ω = 268.^{o}221
Ω = 280.295 }2000.0
q = 3.34281 AU  i = 148.773

SUPERNOVA 2005cz IN NGC 4589 D. C. Leonard, California Institute of Technology, reports that a CCD

spectrum (range 340–900 nm) of SN 2005cz (IAUC 8569), obtained on July 28 UT with the Keck I 10-m telescope (+ LRISp), shows it to be a type-Ib supernova, ∼ 1–2 weeks after maximum light. Although somewhat contaminated by host-galaxy light, the spectrum is quite similar to SN 1998dt at 8 days past maximum (see Matheson et al. 2001, A.J. 121, 1648).

V5116 SAGITTARII R. W. Russell, D. K. Lynch, and R. J. Rudy, The Aerospace Corpo-

ration, report 0.8–2.5-μm spectroscopy of V5116 Sgr on July 15 UT using SpeX at the Infrared Telescope Facility. The nova showed emission lines of H I, He I, C I, N I, Ca II, and O I with FWHM ∼ 2200 km/s. P-Cyg profiles were observed on the He I lines at 1.0830 and 2.0581 μm. There was no evidence of thermal emission from dust.

Additional visual magnitude estimates (cf. IAUC 8561): July 9.892 UT, 9.4 (A. Kammerer, Malsch, Germany); 12.93, 9.3 (J. Carvajal, Madrid, Spain); 30.91, 10.6 (J. G. de S. Aguiar, Campinas, Brazil).

2005 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green